runi Weygandt, Accounting Principles, 12e AND, INC . 170,000
media%2F478%2F4789df8e-3f02-4a2e-938e-c3
Answer #1
Particulars amount
NO OF weighted Avg share 30000
Profit after tax 65000
Preferred dividend 5000
Earning available to equity(65000-5000) 60000
EPS(60000/30000) 2
Dividend per share (26000/30000) . 866666
Payout ratio=DPS/EPS*100 43.33%
Market price of stock 13
PE RATIO=PRICE/EARNINGS=13/2 6.5TIMES
Interest expense 16000
Other expense (105000-24000-16000) 65000

Profit before tax and interest=gross profit - other expense

=170000-65000=105000

Interest earned ratio

=profit before interest and tax/interest expense

=105000/16000=6.5625 times or 6.6times